OpenStudy (anonymous):

A farmer wishes to enclose an 8000 sq meter field and subdivide the field into four rectangular plots, which requires three inner fences each parallel to each other and one of the sides of the field. What should the dimensions of the field be in order to minimize the amount of fencing required?

OpenStudy (anonymous):



OpenStudy (anonymous):

given: xy=8000

OpenStudy (anonymous):

minimize: 5x+2y

OpenStudy (anonymous):

y = 8000 / x y=5x + 16000/x Find the derivative of this equation

OpenStudy (anonymous):

moneybird how did you get those

OpenStudy (anonymous):

y' = 5-16000/x^2 5-16000/x^2 = 0 x = 40sqrt(2) y = 8000/ 4sqrt(2)